OpenStudy (anonymous):

Find all solutions of the following equation. sin x cos x - 2 sin x = 0

OpenStudy (raden):

sin x cos x - 2 sin x = 0 factor out it, get sin x (cos x - 2) = 0 set up each factor be equal zero then solve for x sin x = 0 x = .... ? cosx - 2 = 0 cos x = 2 there is no solution for x

OpenStudy (anonymous):

how did you get to sin x (cosx-2)? what happened to the -2 sin x?

Parth (parthkohli):

He factorized it. For example,\[xy - 2y = y(x - 2)\]

OpenStudy (anonymous):

okay, so if sin x = 0, wouldn't x be equal to 0pie and pie?

OpenStudy (anonymous):

lol pi*